# Density-based clustering algorithm for galaxy group/cluster identification

### v1

@@ Abstract
A direct approach to studying the galaxy-halo connection is the analysis of observed groups and clusters of galaxies that trace the underlying dark matter halos, making identifying galaxy clusters and their associated brightest cluster galaxies (BCGs) crucial. We test and propose a robust density-based clustering algorithm that outperforms the traditional Friends-of-Friends (FoF) algorithm in the currently available galaxy group/cluster catalogs. Our new approach is a modified version of the Ordering Points To Identify the Clustering Structure (OPTICS) algorithm, which accounts for line-of-sight positional uncertainties due to redshift space distortions by incorporating a scaling factor, and is thereby referred to as sOPTICS. When tested on both a galaxy group catalog based on semi-analytic galaxy formation simulations and observational data, our algorithm demonstrated robustness to outliers and relative insensitivity to hyperparameter choices. In total, we compared the results of eight clustering algorithms. The proposed density-based clustering method, sOPTICS, outperforms FoF in accurately identifying giant galaxy clusters and their associated BCGs in various environments with higher purity and recovery rate, also successfully recovering 115 BCGs out of 118 reliable BCGs from a large galaxy sample.
@@
pubyear: 2024pagerange: Density-based clustering algorithm for galaxy group/cluster identificationDensity-based clustering algorithm for galaxy group/cluster identification
@@ 1 Introduction
In addition, central to galaxy clusters are the Brightest Cluster Galaxies (BCGs) located at the bottom of gravitational well within the clusters (Quintana & Lawrie, 1982). The properties of BCGs dictate cluster formation and evolution, where BCG mass growth is closely tied to the hierarchical assembly and dynamical state of the host galaxy cluster (Sohn et al., 2021). What is more distinct from other galaxies is that some of the BCGs show multiple nuclei (e.g. Lauer, 1988; Kluge et al., 2020), making them good systems to study about galactic mergers. A recent study on velocity dispersion profiles of elliptical galaxies also found the majority of the BCGs exhibit flat velocity dispersion profiles (Tian et al., 2021; Duann et al., 2023). A distinct radial acceleration relation (RAR) has even been identified in BCGs, making them essential by posing a significant challenge to the cold dark matter (CDM) paradigm (Tian et al., 2024). However, identifying BCGs can be complex, often requiring a comprehensive survey of galaxies and coherently identifying a pure and complete galaxy clusters catalog first.
@@
However, DBSCAN has limitations, particularly in handling datasets with varying density clusters. Since it relies on a single density threshold to define clusters, DBSCAN can struggle to effectively identify clusters of varying densities. To address these shortcomings, algorithms like Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN; Campello et al., 2015; McInnes et al., 2017) and Ordering Points To Identify the Clustering Structure (OPTICS; Ankerst et al., 1999) have been introduced. In these algorithms, dense points remain the same distance with each other, but noise points are pushed away from any other point. This effectively ‘lowers the sea level’ spreading sparse ‘sea’ points out while leaving ‘land’ untouched. These enhancements make HDBSCAN and OPTICS more suitable for identifying clusters and groups of stars and galaxies that more closely resemble the real distribution and density variations them (Brauer et al., 2022; Fuentes et al., 2017; Oliver et al., 2021).
@@
Therefore, in this work, we conduct comprehensive tests on various clustering methods to explore the possibilities and challenges of identifying galaxy groups and clusters from large galaxy surveys and propose a new algorithm. This paper is organized as follows: Section 2 provides a concise introduction to the clustering algorithms used in this study, detailing the methodology for feature extraction and hyperparameter optimization, including the selection criteria. Section 3 offers a comprehensive evaluation of the effectiveness of group finders using a galaxy catalog derived from simulations. Following this, Section 4 presents additional tests conducted with real-world observations, which include mitigating redshift space distortion using our proposed line-of-sight scaling factor and comparisons with a reliable group catalog. Section 5.1 discusses the strength of our sOPTICS method and its efficiency in identifying BCGs. Finally, Section 6 summarizes our findings and provides a detailed discussion of the results.
@@ 2 Clustering Methodology
Clustering, a technique in machine learning, groups similar data points into clusters and is a powerful tool for identifying galaxy groups in astronomical data. Determining the most advanced clustering algorithm in data science is challenging, as it largely depends on the specific use case and data type. Importantly, a more advanced algorithm is not necessarily a better fit for every dataset or task. This consideration is crucial when employing these algorithms for practical applications, such as extracting galaxy groups and clusters from observations. Hence, an algorithm should be selected based on its suitability for the specific problem. In our research, we have applied eight different clustering algorithms to the 3D spatial positions of simulated galaxies in comoving space. These algorithms include $k$-means (MacQueen, 1967), Gaussian Mixture Models (GMMs) (Dempster et al., 1977), Spectral Clustering (Ng et al., 2001; von Luxburg, 2007), Agglomerative Clustering (Ward, 1963), as well as FoF, DBSCAN, HDBSCAN, and OPTICS, as introduced in Section 1. First, we briefly describe each of these algorithms below.
2.1 Clustering Algorithms
Among all the clustering algorithms considered here, $k$-means stands out for its simplicity of implementation and computational efficiency. The core principle behind $k$-means is to minimize the total intra-cluster variance by constructing clusters and their corresponding centroids. However, traditional $k$-means and its variants are susceptible to local minima in the minimum-sum-of-squares objective function (Bottou & Bengio, 1994). This implies that the clustering result for a given dataset may vary depending on the randomly initialized centroids, impacting the overall process. Consequently, while computationally inexpensive, $k$-means is often utilized as a starting point for computationally heavier algorithms, such as GMMs.
GMMs are generative models capable of representing data with multiple underlying modes or clusters as a linear combination of weighted Gaussian distributions. GMMs are well-suited for soft clustering tasks, where each data point can belong to multiple clusters with varying degrees of membership represented by probabilities. However, GMMs generally require more iterations of the expectation–maximization (EM) algorithm (Dempster et al., 1977) to converge compared to $k$-means, resulting in slower execution times.
Spectral clustering operates under the principle of graph partitioning. It first constructs a similarity graph where nodes represent data points and edges represent similarities between them. Then, it analyzes the eigenvectors (spectrum) of the Laplacian matrix of the Laplacian matrix of this graph to project the data into a lower-dimensional space where traditional clustering algorithms (e.g., $k$-means) can be performed. This approach allows spectral clustering to excel in identifying clusters not linearly separable in the original data space.
Agglomerative clustering employs a bottom-up approach, progressively merging clusters based on their proximity. In this work, we choose the Ward linkage method (Ward, 1963), which minimizes the total within-cluster variance. At each step, the pair of clusters with the minimum between-cluster distance are merged, leading to clusters that are internally coherent and well-separated from each other.
@@ 2.1 Clustering Algorithms
The FoF algorithm is the most popular way of identifying galaxy clusters within a cosmic structure. Given a set of points in space, the FoF algorithm links points within a predetermined distance $l$ to identify interconnected clusters. Two points are considered ‘friends’ (i.e., part of the same cluster) if they are within $l$ of each other. This process iteratively groups points together by linking points to their friends and friends of friends.
@@
The last three algorithms, DBSCAN, HDBSCAN, and OPTICS, are all density-based clustering algorithms. For each point in the dataset, DBSCAN calculates the number of points within a specified radius $\epsilon$. If this number exceeds a minimum number of neighbors $N_{\mathrm{min}}$, the point is classified as a core point, indicating a high-density area surrounding it. The core distance $d_{\mathrm{core}}$ is the minimum radius of a neighborhood around this point that must contain a certain number of other points to qualify the point as a core point. These core points serve as the seeds for cluster growth, as the algorithm iteratively adds directly reachable points (points located within the $\epsilon$-radius of a core point) to their respective clusters. Points not reachable from any core point are labeled as noise.
@@
where $d(P,Q))$ is the Euclidean distance between two points. HDBSCAN then constructs a minimum spanning tree (MST: e.g., Foulds 1991), which connects all data points in a way that the total sum of edge lengths (distances) is minimized. By systematically removing the longest edges from the MST, HDBSCAN creates a dendrogram that reflects the data structure at varying density levels. Each cluster’s stability is calculated as the sum of the excess of density (over a minimum cluster size threshold) for each point within the cluster across the range of distance scales. Finally, HDBSCAN iteratively prunes this dendrogram using the stability criterion, resulting in robust and persistent clusters over a range of densities. (Campello et al., 2013)
Instead of a global $\epsilon$ parameter, OPTICS uses reachability to create an ordered list reflecting the data structure. This ordered list is constructed by iteratively updating the reachability distance for each data point. Starting with an arbitrary point, its reachability distance is calculated relative to its neighbors. This point is added to the list, and the algorithm progresses to the unprocessed point with the smallest reachability distance. This process continues until all points are ordered. This resulting list effectively captures the density-based clustering structure without explicitly assigning points to clusters. Clusters can be extracted by identifying valleys (i.e., low reachability distance regions indicating dense areas) separated by peaks (i.e., high reachability distances marking transitions between clusters or noise). In practice, the $\xi$ parameter defines what constitutes a "steep" decrease or increase: a point is part of a steep downward (or upward) area if its reachability distance is less (or greater) than $\xi$ times the reachability distance of the preceding point in the ordered list.
@@ 2.2 Hyperparameter Optimization
Hyperparameter selection for each of the algorithms in our task requires careful consideration, including the characteristics of the target groups or clusters, data resolution, and noise levels. The following is a breakdown of key hyperparameters for each method.
$k$-means, GMM, Spectral Clustering, and Agglomerative Clustering all share the critical hyperparameter of the number of clusters, $N_{\mathrm{C}}$. Finding the optimal $N_{\mathrm{C}}$ requires careful consideration of the data complexity and desired level of cluster granularity. The FoF method critically relies on the linking length, $l_{\mathrm{linking}}$ to define the scale at which structures are identified.
For DBSCAN, the two primary hyperparameters are $N_{\mathrm{min}}$ and $\epsilon$, which control the degree to which a point is considered a core point. Higher $\epsilon$ values incorporate distant galaxies into clusters, facilitating the identification of larger groups but potentially over-grouping due to projection effects. Conversely, a higher $N_{\mathrm{min}}$ threshold helps identify more significant clusters, reducing the possibility of detecting spurious or minor groupings. HDBSCAN relies on an additional layer of complexity with the min_cluster_size parameter ($M_{\mathrm{min}}$). This parameter sets the minimum number of members required for a grouping to be considered a valid cluster, allowing for the filtering of insignificant or noisy structures. Besides, there is an additional hyperparameter $\alpha$, which relates to the minimum cluster stability required for a cluster to be considered significant. A lower value of $\alpha$ makes it easier for points to be included in a cluster, potentially leading to larger and less dense clusters.
In the case of OPTICS, there are four key hyperparameters, with $N_{\mathrm{min}}$ and $\epsilon$ exerting the most influence. Beyond these two, the $\xi$ parameter plays a crucial role, with lower values leading to the identification of more clusters, even capturing subtle variations in galaxy density. This sensitivity is advantageous for distinguishing closely spaced or subtly different groups. Finally, the minimum number of member galaxies per group, $M_{\mathrm{min}}$, is a straightforward parameter set based on the desired group scale. In the following tests and results, we set $M_{\mathrm{min}}=5$ to focus on groups with at least five members.
All clustering algorithms in this work require a preselected hyperparameter carefully considered for the desired galaxy group scale. Choosing the optimal hyperparameter values involves balancing the preservation of large-scale structures against the fragmentation of real galaxy groups into smaller, potentially insignificant groups. To optimize hyperparameter values, we adopted two classical criteria, purity and completeness, to evaluate the performance of clustering algorithms under different hyperparameter settings. However, comparing results to a simulated group catalog introduces inherent biases. Simulations, while valuable, are not perfect representations of real galaxy groups, and even semi-analytic galaxy group catalogs are usually constructed using FoF clustering algorithms (Onions et al., 2012, also see Section 3.1), introducing bias in the "ground truth" data. Consequently, demanding complete overlap between predicted and simulated groups is unrealistic and unnecessary. Instead, similar to Brauer et al. (2022), we define a broader measure of purity and completeness, incorporating what we term as soft criteria, to assess the performance of the clustering algorithms for a more nuanced evaluation. Under these criteria, a cluster is considered pure if at least two-thirds of its galaxies originated from a single group and complete if it contains at least half the galaxies from that originating group. Building upon the definitions of purity and completeness, we can define the purity rate and recovery rate for all predicted groups relative to the full set of true groups:
@@
When calculating purity and recovery rates, we only compare predicted groups to true groups exceeding the minimum member threshold, $M_{\mathrm{min}}$=5. It is worth noting that, to calculate purity, the traversal list here is all the predicted groups, not the halo IDs in the simulation catalog. Meanwhile, to calculate the recovery rate, the traversal list is all the halos with at least $M_{\mathrm{min}}$ galaxy members in the simulation catalog rather than the predicted groups. Consequently, the purity rate reflects the proportion of predicted groups exclusively containing members from a single true group. It provides confidence for ensuring that the members within a predicted group are genuinely bound together. A high purity rate indicates that the algorithm is effective in correctly grouping members. On the other hand, the recovery rate measures the percentage of true, significant groups successfully identified and reproduced by the algorithm. This ensures the informativeness and reliability of the results for further analysis.

### v2

# sOPTICS: A Modified Density-Based Algorithm for Identifying Galaxy Groups/Clusters and Brightest Cluster Galaxies

@@ Abstract
A direct approach to studying the galaxy-halo connection is to analyze groups and clusters of galaxies that trace the underlying dark matter halos, emphasizing the importance of identifying galaxy clusters and their associated brightest cluster galaxies (BCGs). In this work, we test and propose a robust density-based clustering algorithm that outperforms the traditional Friends-of-Friends (FoF) algorithm in the currently available galaxy group/cluster catalogs. Our new approach is a modified version of the Ordering Points To Identify the Clustering Structure (OPTICS) algorithm, which accounts for line-of-sight positional uncertainties due to redshift space distortions by incorporating a scaling factor, and is thereby referred to as sOPTICS. When tested on both a galaxy group catalog based on semi-analytic galaxy formation simulations and observational data, our algorithm demonstrated robustness to outliers and relative insensitivity to hyperparameter choices. In total, we compared the results of eight clustering algorithms. The proposed density-based clustering method, sOPTICS, outperforms FoF in accurately identifying giant galaxy clusters and their associated BCGs in various environments with higher purity and recovery rate, also successfully recovering 115 BCGs out of 118 reliable BCGs from a large galaxy sample. Furthermore, when applied to an independent observational catalog without extensive re-tuning, sOPTICS maintains high recovery efficiency, confirming its flexibility and effectiveness for large-scale astronomical surveys.
@@
pubyear: 2025pagerange: sOPTICS: A Modified Density-Based Algorithm for Identifying Galaxy Groups/Clusters and Brightest Cluster GalaxiessOPTICS: A Modified Density-Based Algorithm for Identifying Galaxy Groups/Clusters and Brightest Cluster Galaxies
@@ 1 Introduction
In addition, central to galaxy clusters are the Brightest Cluster Galaxies (BCGs) located at the bottom of gravitational well within the clusters (Quintana & Lawrie, 1982). The properties of BCGs dictate cluster formation and evolution, where BCG mass growth is closely tied to the hierarchical assembly and dynamical state of the host galaxy cluster (Sohn et al., 2021). What is more distinct from other galaxies is that some of the BCGs show multiple nuclei (e.g. Lauer, 1988; Kluge et al., 2020), making them good systems to study about galactic mergers. A recent study on velocity dispersion profiles of elliptical galaxies also found the majority of the BCGs exhibit flat velocity dispersion profiles (Tian et al., 2021; Duann et al., 2023). A distinct radial acceleration relation (RAR) has even been identified in BCGs, making them essential to study as they pose a significant challenge to the cold dark matter (CDM) paradigm (Tian et al., 2024). However, identifying BCGs can be complex, often requiring a comprehensive survey of galaxies and coherently identifying a pure and complete galaxy clusters catalog first.
@@
However, DBSCAN has limitations, particularly in handling datasets with clusters of varying densities. Since it relies on a single density threshold to define clusters, DBSCAN can struggle to effectively identify clusters with differing density levels. To address these shortcomings, algorithms such as Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN; Campello et al., 2015; McInnes et al., 2017) and Ordering Points To Identify the Clustering Structure (OPTICS; Ankerst et al., 1999) have been introduced. These algorithms improve upon DBSCAN by adapting to local density variations. In both HDBSCAN and OPTICS, the process can be visualized as ‘lowering the sea level’ in a topographical representation of the dataset: sparse points, which correspond to noise, are gradually pushed away from denser regions (the ‘land’), spreading them out as the sea level drops. Meanwhile, the denser regions, representing clusters, remain largely unchanged, as they stand above the receding sea level. This approach ensures that noise is separated from clusters, allowing for the identification of groups of stars and galaxies that better reflect the real distribution and density variations. (Brauer et al., 2022; Fuentes et al., 2017; Oliver et al., 2021).
@@
Therefore, in this work, we conduct comprehensive tests on various clustering methods to explore the possibilities and challenges of identifying galaxy groups and clusters from large galaxy surveys and propose a new algorithm. This paper is organized as follows: Section 2 provides a concise introduction to the clustering algorithms used in this study, detailing the methodology for feature extraction and hyperparameter optimization, including the selection criteria. Section 3 offers a comprehensive evaluation of the effectiveness of group finders using a galaxy catalog derived from simulations. Following this, Section 4 presents additional tests conducted with real-world observations, which include mitigating redshift space distortion using our proposed line-of-sight scaling factor and comparisons with a reliable group catalog. Section 5 discusses the strength of our sOPTICS method and its efficiency in identifying galaxy groups/clusters and BCGs. Finally, Section 6 summarizes our findings and provides a detailed discussion of the results.
@@ 2 Clustering Methodology
Clustering, a key machine learning technique, groups similar data points into clusters and is a powerful tool for identifying galaxy groups/clusters in astronomical data. Among the available clustering algorithms, the most straightforward and widely used is $k$-means (MacQueen, 1967), often favored for their simplicity and computational efficiency in partitioning data into distinct groups. Beyond this foundational method, more advanced algorithms, such as the Gaussian Mixture Models (GMMs) (Dempster et al., 1977), Spectral Clustering (Ng et al., 2001; von Luxburg, 2007) and Agglomerative Clustering (Ward, 1963), are frequently employed across diverse domains due to their flexibility in identifying more complex or hierarchical data structures. They all share the critical hyperparameter of the number of clusters, $N_{\mathrm{C}}$.
However, notably, a more advanced algorithm may not always be the optimal choice for every dataset or task, as it heavily depends on the specific use case and data type. Therefore, algorithm selection should prioritize suitability for the problem, especially when extracting galaxy groups/clusters from observations. Thus, a comprehensive comparative study testing the suitability of various clustering algorithms is essential to identify the most effective approaches for this task.
2.1 Fiducial Clustering Algorithms
Before presenting the results of applying these algorithms to identify galaxy groups and clusters, we briefly describe the fiducial algorithms employed in our analysis.
2.1.1 Friends-of-Friends (FoF)
@@
In the specific context of identifying galaxy groups or clusters based on their spatial distributions, the Friends-of-Friends (FoF) algorithm stands out as the most commonly used approach. Its popularity arises from its simplicity and its inherent ability to effectively capture the hierarchical and complex characteristics of cosmic structures. Given a set of points in space, the FoF algorithm links points within a predetermined distance $l$ to identify interconnected clusters. Two points are considered ‘friends’ (i.e., part of the same cluster) if they are within $l$ of each other. This process iteratively groups points together by linking points to their friends and friends of friends.
The FoF method critically relies on the linking length, $l$, to define the scale at which structures are identified. That is, smaller values of $l$ lead to the identification of smaller, tighter clusters by focusing on locally dense regions, but may fail to connect broader structures. Conversely, larger values of $l$ allow for the detection of more extended, diffuse clusters by linking points over greater distances. However, overly large $l$ values risk merging distinct structures into a single cluster, reducing the purity of the results. As a result, careful tuning of $l$ is crucial to balance the trade-off between identifying small-scale structures and preserving the integrity of larger-scale groupings.
2.1.2 DBSCAN
As mentioned in Section 1, density-based clustering algorithms such as DBSCAN, HDBSCAN, and OPTICS have been relatively underutilized or insufficiently explored in the context of identifying galaxy groups or clusters. However, these methods hold significant potential due to their capability to identify clusters of arbitrary shapes and effectively handle noise in the data. In this work, we focus on an in-depth investigation of these algorithms, examining their suitability and performance for extracting galaxy groups and clusters from observational datasets.
@@ 2.1.2 DBSCAN
For each point in the dataset, DBSCAN calculates the number of points within a specified radius $\epsilon$. If this number exceeds a minimum number of neighbors $N_{\mathrm{min}}$, the point is classified as a core point, indicating a high-density area surrounding it. The core distance, $d_{\mathrm{core}}(P,N_{\mathrm{min}})$, is the distance from a point P to its $N_{\mathrm{min}}$-th most nearest neighbor and if $d_{\mathrm{core}}(P,N_{\mathrm{min}})\leq\epsilon$ then P is a core-point. These core points serve as the seeds for cluster growth, as the algorithm iteratively adds directly reachable points (points located within the $\epsilon$-radius of a core point) to their respective clusters. Points not reachable from any core point are labeled as noise.
For DBSCAN, the two primary hyperparameters are $N_{\mathrm{min}}$ and $\epsilon$, which together determine how a point qualifies as a core point. Larger $\epsilon$ values allow for the inclusion of more distant galaxies within clusters, enabling the identification of larger, more extended groups but increasing the risk of over-grouping due to projection effects. Conversely, higher $N_{\mathrm{min}}$ thresholds help identify more substantial clusters by focusing on denser regions, reducing the likelihood of detecting spurious or minor groupings. Specifically in this work, we introduce an additional layer of complexity with the min_member parameter ($M_{\mathrm{min}}$) to DBSCAN. This parameter specifies the minimum number of members required for a grouping to qualify as a valid cluster, enabling the algorithm to filter out insignificant or noisy structures effectively. The $M_{\mathrm{min}}$ parameter is also applied in HDBSCAN and OPTICS in the subsequent discussion.
2.1.3 HDBSCAN
@@
$d_{\mathrm{reach}}(P,Q)=\max\{d_{\mathrm{core}}(P,N_{\mathrm{min}}),d_{\mathrm{core}}(Q,N_{\mathrm{min}}),d(P,Q)\}\,,$ (1)
where $d(P,Q))$ is the Euclidean distance between two points. HDBSCAN then constructs a minimum spanning tree (MST: e.g., Foulds 1992), which connects all data points in a way that the total sum of edge lengths (distances) is minimized. By systematically removing the longest edges from the MST, HDBSCAN creates a dendrogram that reflects the data structure at varying density levels. Each cluster’s stability is calculated as the sum of the excess of density (over a minimum cluster size threshold) for each point within the cluster across the range of distance scales. Finally, HDBSCAN iteratively prunes this dendrogram using the stability criterion, resulting in robust and persistent clusters over a range of densities. (Campello et al., 2013)
Comparing to DBSCAN, HDBSCAN incorporates another key hyperparameter $\alpha$, which governs the minimum stability a cluster must achieve to be considered significant. Lower values of $\alpha$ reduce the stability threshold, making it easier for points to be included in a cluster. While this can lead to larger and less dense clusters, it may also increase the likelihood of capturing subtle structures at the cost of potential over-grouping.
2.1.4 OPTICS
Instead of relying solely on a global $\epsilon$ parameter, OPTICS uses reachability to create an ordered list that reflects the data structure. The reachability distance in OPTIC is defined as:
@@ 2.1.4 OPTICS
which differs slightly from HDBSCAN’s approach, as it is specifically designed to construct the ordered list. The ordered list is built by iteratively calculating and updating the reachability distance for each data point. Starting with an arbitrary point, its reachability distance is calculated relative to its neighbors. This point is added to the list, and the algorithm progresses to the unprocessed point with the smallest reachability distance. This process continues until all points are ordered. The resulting list encapsulates the density-based clustering structure without explicitly assigning points to clusters. Clusters can be then extracted by identifying valleys (low reachability distances, representing dense areas) separated by peaks (high reachability distances, marking transitions between clusters or noise) in the reachability plot. The $\xi$ parameter defines what constitutes a “steep” change in the reachability plot: it is used to detect sharp transitions (e.g., steep downward or upward slopes) by comparing the relative change in reachability distance between consecutive points in the ordered list. Specifically, $\xi$ marks the boundaries of these changes, identifying where clusters start or end, rather than marking all points within a region of steep gradient. This approach ensures that the clustering structure is captured at multiple density scales while avoiding over-segmentation of data.
Similar to DBSCAN, the parameter $N_{\mathrm{min}}$, which defines the minimum number of neighbors required to consider a point as a core point, controls the algorithm’s sensitivity to local density variations. Smaller values of $N_{\mathrm{min}}$ allow for the identification of smaller, more localized clusters, but may also increase the detection of noise and spurious groupings. The $\epsilon$ parameter determines the maximum distance for evaluating reachability, and while it does not directly dictate cluster boundaries as in DBSCAN, it sets an upper limit for defining local neighborhoods. Finally, the $\xi$ parameter, which identifies steep changes in the reachability plot, governs the resolution of cluster extraction. Lower $\xi$ values detect finer density variations, potentially identifying smaller or closely spaced clusters, whereas higher $\xi$ values focus on broader, more prominent clustering structures. Together, these parameters provide OPTICS with the flexibility to adapt to various density scales and cluster complexities.
@@ 2.2 Hyperparameter Optimization
All clustering algorithms in this work require a preselected hyperparameter carefully considered for the desired galaxy group scale. Choosing the optimal hyperparameter values involves balancing the preservation of large-scale structures against the fragmentation of real galaxy groups into smaller, potentially insignificant groups. To optimize hyperparameter values, we adopt two classical criteria, purity and completeness, to evaluate the performance of clustering algorithms under different hyperparameter settings. However, comparing results to a simulated group catalog introduces inherent biases. Simulations, while valuable, are not perfect representations of real galaxy groups, and even semi-analytic galaxy group catalogs are usually constructed using FoF clustering algorithms (Onions et al., 2012, also see Section 3.1), introducing bias in the "ground truth" data. Consequently, demanding complete overlap between predicted and simulated groups is unrealistic and unnecessary. Instead, similar to Brauer et al. (2022), we define a broader measure of purity and completeness, incorporating what we term as soft criteria, to assess the performance of the clustering algorithms for a more nuanced evaluation. Under these criteria, a cluster is considered pure if at least two-thirds of its galaxies originated from a single group and complete if it contains at least half the galaxies from that originating group. Building upon the definitions of purity and completeness, we can define the purity rate and recovery rate for all predicted groups relative to the full set of true groups:
@@
When calculating purity and recovery rates, we only compare predicted groups to true groups exceeding the minimum member threshold, $M_{\mathrm{min}}$=5. It is worth noting that, to calculate purity, the traversal list is defined as all the predicted groups in the clustering results, not the true groups in the simulation catalog. Specifically, for calculating the purity rate, we evaluate each predicted group to determine whether its members correspond to a single true group in the simulation catalog, rather than starting from the simulation catalog to find the corresponding predicted group. Conversely, for calculating the recovery rate, the traversal list consists of all halos with at least $M_{\mathrm{min}}$ galaxy members in the simulation catalog, rather than the predicted groups. As a result, the purity rate reflects the proportion of predicted groups whose members are exclusively from a single true group, providing confidence that the algorithm correctly groups members together. A high purity rate indicates the algorithm’s effectiveness in identifying true groupings. Consequently, the purity rate reflects the proportion of predicted groups exclusively containing members from a single true group. It provides confidence for ensuring that the members within a predicted group are genuinely bound together. A high purity rate indicates that the algorithm is effective in correctly grouping members. On the other hand, the recovery rate measures the percentage of true, significant groups successfully identified and reproduced by the algorithm. This ensures the informativeness and reliability of the results for further analysis.